Find the sum of squares of all numbers x such that both expressions 9x - x^2 and x + (1/x) are integer numbers.
1
Expert's answer
2020-11-09T20:33:21-0500

Let "9x-x^2=m" and "x+\frac{1}{x}=n" , where "m" and "n" are integer numbers.

If "x""\ 0" , then

"\begin{cases}
9x-x^2=m
\\
x+\frac{1}{x}=n
\end{cases}
\quad \Leftrightarrow 
\quad
 \begin{cases}
x^2-9x+m=0
\\
x^2+1=nx 
\end{cases}
(1)
\quad \Leftrightarrow 
\quad 
\begin{cases}-9x+m-1=-nx
\\
x^2-nx+1=0
\end{cases}"

"x(9-n)=m-1"

If "n=9" , then "m=1" and system of equations "(1)" is equivalent to "x^2-9x+1=0" .

"x_{1,2}=\frac{9\pm \sqrt{81-4}}{2}=\frac{9\pm\sqrt{77}}{2}" , and "(x_1)^2+(x_2)^2=\frac{81+18\sqrt{77}+77}{4}+ \frac{81-18\sqrt{77}+77}{4}=79"

If "n""9" , then "x=\frac{m-1}{9-n}" is a rational number.

"x" is one of the roots of "x^2-nx+1=0" . The roots are "x_{1,2}=\frac{n\pm\sqrt{n^2-4}}{2}" .

We know, that "x" is rational, therefore "\sqrt{n^2-4}" is integer, because square root of integer number is either integer or irrational. If it was irrational, then fraction "\frac{n\pm\sqrt{n^2-4}}{2}" wouldn’t be a rational number.

So, "n^2-4=k^2" , where "k" is integer.

"n^2-k^2=4\quad \Leftrightarrow \quad (n-k)(n+k)=4"

"n-k" and "n+k" are both even numbers (if they are odd, then their product is odd; if one number is odd and another one is even, then their sum is odd, but it is equal to "2n" )

We have two cases:

1) "\begin{cases}
n-k=2
\\
n+k=2
\end{cases}" and then "n=2", "k=0" .

2) "\begin{cases}
n-k=-2
\\
n+k=-2
\end{cases}" and then "n=-2", "k=0."

If "n=2" : "x^2-2x+1=(x-1)^2=0" and "x=1" . If this case "9x-x^2" and "x+\frac{1}{x}" are integers.

If "n=-2" : "x^2+2x+1=(x+1)^2=0" and "x=-1" . If this case "9x-x^2" and "x+\frac{1}{x}" are integers.


So, the sum of squares is "79+1+1=81"


Answer: the sum of squares of all numbers x such that both expressions "9x-x^2" and "x+\frac{1}{x}" are integer numbers is equal to 81.